# ISRE-$0$-$0$.00

### v2

@@ 6 Results
Following the estimation procedure described earlier, the results are summarized in Table 4. Firstly, in response to RQ2, we document the existence of a utility divide in real-world contexts, demonstrating that different demographic groups derive varying levels of utility from ChatGPT compared to traditional search engines. Specifically, with regard to age and gender, our analysis reveals that younger users and males derive significantly greater benefit from ChatGPT, reporting higher utility levels ($\alpha_{2}=-0.021$ and $\alpha_{3}=0.562$). These findings are consistent with the extant literature on the digital divide, which has repeatedly shown that younger and male users tend to exhibit higher rates of ICT adoption and usage (Yang and Zhang 2023, Hargittai 2002, Jackson et al. 2001). Regarding occupational differences, our model suggests that individuals employed in IT-related sectors experience enhanced utility from ChatGPT ($\alpha_{5}=0.507$), likely due to their higher levels of digital literacy. This observation aligns with previous studies on the digital divide, which have emphasized that individuals with more advanced computer skills are better positioned to utilize ICT effectively (Hargittai 2003, Van Dijk and Hacker 2003). Moreover, recent research on AI adoption, such as that by Alekseeva et al. (2020), has identified a positive relationship between AI adoption and increased firm growth and investment, particularly in IT-intensive industries.
@@
$\lambda$ Latent class probability parameter --0.384 $(0.077)^{***}$
@@
$c$ Mean representative utility of traditional search engines 1.411 $(0.013)^{***}$
$\alpha_{0}$ Intercept of ChatGPT’s representative utility --1.560 $(0.024)^{***}$
@@
$log(\Delta\alpha_{0})$ Difference in ChatGPT’s representative utility between two latent classes 0.976 $(0.005)^{***}$
$\alpha_{1}(HighEdu)$ Coefficient of high education levels on ChatGPT’s representative utility --0.468 $(0.011)^{***}$
@@
$\alpha_{2}(Age)$ Coefficient of ages on ChatGPT’s representative utility --0.021 $(0.001)^{***}$
$\alpha_{3}(Male)$ Coefficient of males on ChatGPT’s representative utility 0.562 $(0.016)^{***}$
$\alpha_{4}(White)$ Coefficient of race (white vs non-white) on ChatGPT’s representative utility --0.208 $(0.011)^{***}$
@@
$\alpha_{5}(IT)$ Coefficient of IT practitioners on ChatGPT’s representative utility 0.507 $(0.021)^{***}$
$\sigma_{n}^{2}$ News signal variance 4.842 $(0.028)^{***}$
$\gamma_{0}$ Intercept of usage signal variance 4.900 $(0.153)^{***}$
@@
$\Delta\gamma_{0}$ Difference in usage signal variance between two latent classes 2.031 $(0.011)^{***}$
$\gamma_{1}(HighEdu)$ Coefficient of high education levels on usage signal variance --0.256 $(0.036)^{***}$
@@
$\gamma_{2}(Age)$ Coefficient of ages on usage signal variance 0.029 $(0.005)^{***}$
$\gamma_{3}(Male)$ Coefficient of males on usage signal variance --0.522 $(0.030)^{***}$
$\gamma_{4}(White)$ Coefficient of white people on usage signal variance --0.481 $(0.129)^{***}$
@@
$\gamma_{5}(IT)$ Coefficient of IT practitioners on usage signal variance --0.456 $(0.045)^{***}$
Note: Standard errors in parentheses; ∗∗∗ p$<$0.01, ∗∗ p$<$0.05, p$<$0.1.
While prior research has investigated adoption behavior in real-world settings and evaluated outcome differences, such as productivity, in laboratory settings, the underlying “unobservable” utility has not been directly inferred from real-world data. Our results contribute to the existing literature by explicitly demonstrating the existence of a utility divide based on demographic attributes.
Moreover, while our findings confirm some of literature’s observations regarding observable behavior, we also present evidence that challenges established notions within the digital divide literature. Most notably, concerning education—traditionally considered one of the most significant predictors of the digital divide (Autor et al. 1998)—we observe that users with lower educational attainment experience substantially higher relative utility from ChatGPT ($\alpha_{1}=-0.468$). This result contradicts prior research suggesting that higher education correlates with improved outcomes in ICT adoption (Scheerder et al. 2017). By disproportionately benefiting users with lower educational attainment, generative AI has the potential to mitigate the outcome divide stemming from educational disparities. Although it is empirically infeasible to test the contributing factors of this phenomenon directly, we provide some intuition and call for further research to gather additional data and study the mechanisms. For example, individuals with different educational backgrounds might possess different needs for online resources, and the needs of less-educated users may align more closely with ChatGPT’s capabilities compared to traditional search engines. Conversely, queries from more educated users may be too complex for direct interaction with ChatGPT, making traditional search engines more suitable. Furthermore, less-educated users may have greater difficulty navigating the results from traditional search engines. Since ChatGPT provides answers in natural language that are easier to understand, less-educated users may derive relatively more benefits from interacting with it, whereas this advantage may be less significant for highly educated individuals.
Furthermore, we also find that white users tend to gain comparatively less utility from ChatGPT ($\alpha_{4}=-0.208$), which contradicts previous literature typically finding that white users have greater access to and derive more benefits from digital technologies (Fairlie 2004). This phenomenon may be attributed to the lower language barriers experienced by white users, leading to relatively diminished marginal utility from ChatGPT compared to traditional search engines.
@@
In addition to the divide in actual utility, we observe a significant deviation of $v_{i}$ from $v_{0}$, where $v_{0}$ is normalized to zero. For example, consider a 25-year-old female, non-white user with low educational attainment, not employed in the IT industry, and belonging to latent class 2. For this individual, the actual representative utility $v_{i}$ is underestimated by $v_{0}-(\alpha_{0}+\Delta\alpha+25\cdot\alpha_{2})=-0.569$. On average, the initial belief regarding ChatGPT’s representative utility is approximately 70% lower than its true representative utility for users in latent class 2. Consequently, without an effective learning process, these users may experience substantial welfare losses in situations where ChatGPT would be more suitable than traditional search engines. Conversely, some individuals overestimate their utility gain from ChatGPT, which may lead them to use ChatGPT when traditional search engines would be more appropriate. Given the pronounced discrepancy between initial beliefs and actual utility, it is crucial for individuals to learn and adjust their perceptions to enable more informed decision-making.
However, the ability to update these beliefs may vary significantly among individuals with different demographic attributes, as highlighted by RQ1 (learning divide). By mathematical definition, the larger the signal variance $\delta_{is}^{2}$, the slower the learning process. Our estimation results substantiate the presence of a significant learning divide. Interestingly, while less-educated and non-white individuals generally benefit more from ChatGPT per use in terms of $v_{i}$, their learning speed is significantly slower, as indicated by the signal variance parameters ($\gamma_{1}=-0.256$ and $\gamma_{4}=-0.481$). Thus, without targeted interventions, lower-educated users may adopt ChatGPT at a slower rate and are more susceptible to extreme belief updates, potentially resulting in suboptimal usage patterns. These heterogeneities help reconcile the apparent contradiction between the utility gains per usage and the observed adoption behavior from the digital divide literature: although less-educated and non-white users exhibit higher utility gains per interaction, their slower learning impedes efficient adoption of ChatGPT. Furthermore, our findings indicate that younger and male users are quicker to perceive the true benefits of ChatGPT ($\gamma_{2}=0.029$ and $\gamma_{3}=-0.522$). Likewise, individuals employed in IT-related industries demonstrate faster learning rates ($\gamma_{5}=-0.456$). Across these dimensions, the learning divide mirrors the utility divide, thereby exacerbating the disparity between different demographic groups.
@@
Finally, it is important to note that our results remain robust when subjected to: (i) alternative value of the prior belief variance (hyperparameter); (ii) the inclusion of different news outlets retrieved in October 2024; and (iii) re-estimating the results after the exclusion of outliers. These robustness checks further reinforce the validity of our findings.
@@ 7 Policy Simulation
First, to demonstrate the importance of learning in the adoption process, we simulate the belief sequence of a typical fast learner. Based on model estimates, we consider a 28-year-old [^5] The 25th percentile of age is used to represent younger individuals. male IT practitioner with a high level of education from latent class 2. This individual faces the choice between ChatGPT and traditional search engines 18 times [^6] The mean of daily activities is used to represent typical behavior. a day over 1000 days. The red dots in Figure 3 illustrate the simulated trajectory of his belief over time. We observe that the user stops using ChatGPT and updating his belief after receiving relatively negative signals during the initial interactions. Despite substantial uncertainty about the true utility of ChatGPT, the user finds it unrewarding to continue using it and instead sticks with search engines. Consequently, he misses the opportunity to update his belief and potentially benefit from ChatGPT. In other words, a few negative experiences can sometimes lead to a persistent negative belief, even for fast learners. We refer to this phenomenon as a belief trap, where users are trapped by their negative beliefs, cease to engage with new technology, and struggle to escape without external intervention.
Given the significant implications of the belief trap, we investigate how the learning divide influences the likelihood of becoming trapped. In addition to the fast learner, for comparison, we consider a typical slow learner: a 48-year-old [^7] We use the 75th percentile of age to represent relatively elder individuals. female, non-white, non-IT practitioner with a low level of education. To control for the impact of unobservables, we also assign this individual to latent class 2. We mathematically define “being trapped” as a user whose current probability of using ChatGPT, based on their present beliefs, is below $1\%$ of the optimal probability determined by the true mean utility $v_{i}$ after one year [^8] The relative relationships depicted in the subsequent Figure 3 remain consistent across different time horizons; however, we present the snapshot trapping probability after a one-year period only for illustrative purposes.. Accordingly, we simulate 10,000 trajectories for each user and plot the probability of becoming trapped. The first two error bars in Figure 3 depict the simulated probabilities of being trapped for the fast learner and the slow learner, respectively, along with 95% confidence intervals calculated using 1,000 bootstrap samples. The slow learner exhibits a significantly higher probability of being trapped, likely due to greater uncertainty in the learning process. To control for the possibility that this difference is driven by heterogeneous utilities—since higher utilities facilitate adoption and thus belief updates—we further simulate a scenario where the slow learner’s true utility is set equal to that of the fast learner. As shown in the third bar of Figure 3, even after controlling for utility, slow learners remain more likely to become trapped. This persistent difference in trapping probability underscores the impact of the learning divide on dynamic adoption behaviors and welfare.
Due to the substantial impact of the learning divide, mitigating this divide and associated consequences becomes crucial. Our simulation results suggest that offering free training could be an effective solution. In Figure 3, the blue dots represent the updated trajectory for the fast learner. By providing free training at the outset and ensuring that he uses ChatGPT 200 times before Day 00, we observe that he is no longer trapped; his belief is frequently updated and approaches his true representative utility of ChatGPT, $0.90$. These findings demonstrate that early training can help users overcome the belief trap. Therefore, offering training to slow learners may help mitigate the learning divide. Furthermore, the last two error bars in Figure 3 illustrate the simulated probability of the slow learner being trapped after 100 and 150 times of interactions, respectively. The results indicate that a training program involving 100–150 uses of ChatGPT enables slow learners to achieve trapping probabilities similar to fast learners, suggesting that this approach can effectively mitigate the learning divide in practice.
Figure 2: Belief Change Over Time
Figure 3: Simulated Probability of Being Trapped
8 Conclusion
Inspired by prior literature on the digital divide, our study introduces two novel concepts to better understand the generative AI divide. First, the learning divide, a critical facet of the capability divide, highlights disparities in individuals’ abilities to effectively update their perceived utility of generative AI through repeated interactions and access to information resources. Second, the utility divide, a form of outcome divide, characterizes differences in the utility derived from each usage of ChatGPT among individuals with different demographics. Utilizing a six-month clickstream dataset, we build and estimate a Bayesian learning model to examine how different demographic groups learn about and derive utility from ChatGPT.
Our analyses reveal significant learning and utility divides across demographics. Males, younger individuals, and those with IT backgrounds not only derive higher utility per use but also update their beliefs about GenAI more rapidly. Conversely, individuals from minority racial groups or with lower educational attainment update their beliefs more slowly, even though they may experience higher utility gains per use. Building on these empirical insights, we conduct simulations using the estimated parameters to model belief sequences over time. Notably, we identify a concerning phenomenon—the belief trap—where users who underestimate generative AI’s utility may abandon the technology, foregoing opportunities for welfare enhancement. The simulations further demonstrate how the learning divide leads to significant differences in the probability of being trapped, thereby contributing to broader outcome divides in adoption dynamics. Finally, we employ counterfactual simulations to show that interventions such as training programs can mitigate these disparities by enhancing exposure and improving belief updating processes.
This study makes several contributions to the literature on the digital divide, AI adoption, and Bayesian learning. First, we develop a novel framework for understanding the AI adoption process in contexts characterized by information uncertainty and dynamic learning. Unlike previous research that often focuses on access and initial usage, our model emphasizes the ongoing learning dynamics that shape user interactions with AI technologies over time. Second, we introduce the concepts of the learning divide, utility divide, and belief trap to elucidate the mechanisms behind differential AI adoption and outcomes. Third, we highlight how AI technologies, without proper intervention strategies, risk amplifying existing social inequalities, emphasizing the critical importance of inclusive design and deployment practices. This highlights the importance of considering equity and inclusion in the design and deployment of AI systems.
Beyond theoretical contributions, our findings have practical implications for governments, NGOs, AI providers, and companies. Governments and NGOs should promote users’ capabilities in using AI through educational programs to help underprivileged users overcome information asymmetry. AI platforms, such as Hugging Face, could implement inclusivity-focused ELO rankings, create inclusive benchmarks, and introduce badges to motivate developers towards inclusivity. Companies that provide AI or integrate AI into their products should prioritize inclusive product design and offer training programs for disadvantaged groups. By addressing the utility divide, learning divide, and belief trap, these stakeholders can promote equitable AI adoption and ensure that the benefits of AI technologies are broadly shared across society.

### v3

@@ 7 Results
Following the estimation procedure described earlier, we summarize the results in Table 2. First, we document a utility divide in a real-world setting, showing that individuals with different social characteristics derive different levels of utility from ChatGPT relative to traditional search engines. Regarding age and gender, we find that younger users and males benefit significantly more from ChatGPT, as reflected in higher estimated utility ($\alpha_{2}=-1.168$, $\alpha_{3}=-0.886$, and $\alpha_{4}=1.403$). Notably, the utility difference between the young group (18 ~ 34) and the middle-aged group (35 ~ 47) is larger than the difference between the middle-aged group and the older group (47+). To further assess the age pattern, we estimate an alternative specification that includes continuous $Age$ and $Age^{2}$ and report the results in Appendix T. The small but positive coefficient on $Age^{2}$ suggests that as age increases, relative utility declines at a slightly decreasing rate. These findings are consistent with the theoretical argument that younger and male users tend to develop higher AI literacy through day-to-day use, and therefore derive greater utility from generative AI relative to traditional search engines (Yan et al. 2024, Robinson and others 2020, Scheerder et al. 2017).
Further, our model indicates that individuals with higher levels of LLM occupational exposure derive greater utility from ChatGPT ($\alpha_{6}=1.058$). This pattern is consistent with prior digital divide research showing that individuals with more advanced computer skills are better positioned to use ICT effectively (Hargittai 2003, Van Dijk and Hacker 2003). In the LLM context, higher $LLMOE$ typically implies that users’ job tasks are more closely aligned with LLM functionalities and that they have more opportunities to incorporate ChatGPT into routine workflows, thereby realizing higher utility (Eloundou et al. 2024).
@@
$\lambda$ Latent class probability parameter -1.336 $(0.071)^{***}$
$c$ Mean representative utility of traditional search engines 0.627 $(0.008)^{***}$
@@
$\Delta c$ Difference in mean representative utility of traditional search engines between two latent classes 0.401 $(0.017)^{***}$
$\alpha_{0}$ Intercept of ChatGPT’s representative utility 0.402 $(0.055)^{***}$
@@
$log(\Delta\alpha_{0})$ Difference in ChatGPT’s representative utility between two latent classes 0.537 $(0.021)^{***}$
$\alpha_{1}(College)$ Coefficient of college education on ChatGPT’s representative utility -0.539 $(0.015)^{***}$
@@
$\alpha_{2}(AgeHigh)$ Coefficient of high age group on ChatGPT’s representative utility -1.168 $(0.022)^{***}$
@@
$\alpha_{3}(AgeMid)$ Coefficient of middle age group on ChatGPT’s representative utility -0.886 $(0.019)^{***}$
$\alpha_{4}(Male)$ Coefficient of males on ChatGPT’s representative utility 1.403 $(0.017)^{***}$
$\alpha_{5}(White)$ Coefficient of race (white vs non-white) on ChatGPT’s representative utility -0.900 $(0.011)^{***}$
$\alpha_{6}(LLMOE)$ Coefficient of LLM occupational exposure on ChatGPT’s representative utility 1.058 $(0.028)^{***}$
$\alpha_{7}(English)$ Coefficient of English literacy on ChatGPT’s representative utility -1.107 $(0.057)^{***}$
$\sigma_{n}^{2}$ News signal variance 5.777 $(0.024)^{***}$
$\gamma_{0}$ Intercept of usage signal variance 5.051 $(0.061)^{***}$
@@
$\Delta\gamma_{0}$ Difference in usage signal variance between two latent classes 3.748 $(0.022)^{***}$
$\gamma_{1}(College)$ Coefficient of college education on usage signal variance -0.272 $(0.017)^{***}$
@@
$\gamma_{2}(AgeHigh)$ Coefficient of high age group on usage signal variance 0.288 $(0.021)^{***}$
@@
$\gamma_{3}(AgeMid)$ Coefficient of middle age group on usage signal variance 0.175 $(0.020)^{***}$
$\gamma_{4}(Male)$ Coefficient of males on usage signal va riance -0.193 $(0.017)^{***}$
$\gamma_{5}(White)$ Coefficient of white people on usage signal variance -0.712 $(0.018)^{***}$
$\gamma_{6}(LLMOE)$ Coefficient of LLM occupational exposure on usage signal variance -0.925 $(0.079)^{***}$
$\gamma_{7}(English)$ Coefficient of English literacy on usage signal variance -0.757 $(0.046)^{***}$
Note: Standard errors in parentheses; ∗∗∗ p<<0.01, ∗∗ p<<0.05, p<<0.1.
While our findings corroborate prior observations for several social characteristics, they also provide evidence that challenges established views about who benefits more in the digital divide literature. Most notably, with respect to education, one of the most prominent predictors of the digital divide (Autor et al. 1998), we find that users without any college education obtain substantially higher relative utility from ChatGPT ($\alpha_{1}=-0.539$). This result appears to contrast with prior research suggesting that education is positively associated with ICT adoption outcomes (Scheerder et al. 2017). However, it can be reconciled by noting that users with some college education may have higher overall literacy and thus can leverage both ChatGPT and traditional search engines more effectively. Because ChatGPT provides natural-language answers that can be easier to interpret, it may function as an “equalizer” that users with and without college education can both utilize. As a result, users without college education may benefit more in relative terms because their baseline utility from traditional search engines is lower.
Furthermore, we find that white users obtain comparatively lower utility from ChatGPT ($\alpha_{5}=-0.900$), which contrasts with prior work that typically finds white users have greater access to and derive more benefits from digital technologies (Fairlie 2004). In addition, users with lower English literacy gain significantly more utility from ChatGPT relative to search engines ($\alpha_{7}=-1.107$). These results suggest another channel through which cultural and language backgrounds shape generative AI use: although the underlying LLM is trained disproportionately on English-language corpora and may align more closely with Western cultural contexts (Agarwal et al. 2025, Martínez-Murillo et al. 2025, Qin et al. 2025), its natural-language interaction mode can also help users from other backgrounds articulate their needs in detail and obtain information that may be harder to navigate through traditional search engines (Xu et al. 2023). As a result, non-white users and those with lower English literacy may experience larger relative gains from ChatGPT compared with traditional search engines.
@@
In addition to the divide in actual utility, we observe a significant deviation of $v_{i}$ from $v_{0}$, where $v_{0}$ is normalized to zero. For example, consider a female user younger than 35 who is non-white, did not receive any college education, has zero English literacy and no LLM occupational exposure, and belongs to latent class 2. For this individual, the actual representative utility $v_{i}$ is underestimated by $(\alpha_{0}+\Delta\alpha)-v_{0}=2.113$. Consequently, without an effective learning process, these users may experience substantial welfare losses due to failure in choosing the more suitable tools. Given the pronounced discrepancy between initial beliefs and actual utility, it is crucial for individuals to learn and adjust their perceptions to enable more informed decision-making.
However, individuals’ ability to update these beliefs may vary substantially across social characteristics (learning divide), and the groups that learn faster may not be the same as those that experience higher per-use utility gains. By definition, a larger signal variance implies slower learning. This divergence is already evident across the latent classes: holding observable social characteristics constant, a user in latent class 2 obtains a larger gain from ChatGPT ($\Delta\alpha_{0}-\Delta c=\exp(0.537)-0.401=1.310$) but learns more slowly ($\Delta\gamma_{0}=3.748$) than a comparable user in latent class 1.
Interestingly, while non-white individuals, users with lower English literacy, and users without college education tend to benefit more from ChatGPT on a per-use basis, they learn significantly more slowly, as reflected in the signal-variance parameters ($\gamma_{1}=-0.272$, $\gamma_{5}=-0.712$, and $\gamma_{7}=-0.757$). Thus, absent targeted interventions, these users may adopt ChatGPT more slowly and may be more prone to extreme belief updates, potentially leading to suboptimal usage patterns. These heterogeneities help reconcile the apparent tension between per-use utility gains and adoption patterns documented in the digital divide literature (Scheerder et al. 2017, Fairlie 2004): although these users obtain higher utility from each interaction, slower learning can impede efficient adoption of ChatGPT.
In contrast, younger and male users update toward the true benefits of ChatGPT more quickly ($\gamma_{2}=0.288$, $\gamma_{3}=0.175$, and $\gamma_{4}=-0.193$). Consistent with the utility estimates, the relationship between learning speed and age is also nonlinear, with learning speed declining at a decreasing rate as age increases. Likewise, we find that individuals with lower occupational exposure to LLMs exhibit faster learning rates ($\gamma_{6}=-0.925$). Across these dimensions, the learning divide partly mirrors the utility divide, thereby amplifying disparities across demographic groups.
@@
Finally, it is important to note that our results remain robust when subjected to: (i) restricting the sample to observations prior to the GPT-4 launch; (ii) alternative value of the prior belief variance (hyperparameter); (iii) the inclusion of different news outlets retrieved in October 2024; (iv) re-estimating the results after the exclusion of outliers; and (v) excluding search queries that are not substitutable by ChatGPT. The results can be found in Appendix S, Appendix U, Appendix V, Appendix W, and Appendix X respectively. These robustness checks further reinforce the validity of our findings.
@@ 8 Policy Simulation
First, to illustrate the role of learning in the adoption process, we simulate the belief trajectory of a typical slow learner. Specifically, we consider a 35-to-47-year-old female non-white user with no English literacy, no college education, and no LLM occupational exposure, who belongs to latent class 2. This individual faces the choice between ChatGPT and traditional search engines 16 times per day [^6] We use the sample mean of daily activities to represent typical behavior. over a horizon of 1,000 days. The simulated belief path is reported in Appendix Y. Despite substantial uncertainty about ChatGPT’s true utility, the user quickly finds continued use unrewarding and instead relies on search engines. As a result, she forgoes opportunities to update her beliefs through further interactions and to potentially benefit from ChatGPT. More generally, a small number of negative experiences can generate persistently pessimistic beliefs. We refer to this phenomenon as a belief trap, in which users become anchored by negative beliefs, discontinue engagement with the new technology, and are unlikely to escape without external intervention.
Given the important implications of the belief trap, we next examine how the learning divide affects the likelihood of becoming trapped. In addition to the slow learner, we consider a representative fast learner for comparison: a male user younger than 35 who is white, has a college education, and has the highest English literacy and LLM occupational exposure. To control for unobserved heterogeneity, we assign this individual to latent class 2 as well. We define “being trapped” as follows: after one year, a user is classified as trapped if their current probability of using ChatGPT, based on their prevailing beliefs, falls below $1\%$ of the optimal probability implied by the true mean utility. We then simulate 10,000 belief trajectories for each user and compute the probability of becoming trapped; the resulting trap rates are reported in Appendix Y. Our results show that the slow learner is substantially more likely to become trapped, plausibly because greater uncertainty slows learning and makes early negative signals more consequential. To rule out the possibility that this gap is driven by heterogeneous utilities, we conduct an additional counterfactual exercise. Because higher utility increases adoption and thus creates more opportunities for belief updating, we set the slow learner’s true utility equal to that of the fast learner and re-run the simulations in a counterfactual scenario. Even under this utility-equalized scenario, slow learners remain more likely to become trapped. This persistent gap in trapping probabilities highlights the role of the learning divide in shaping dynamic adoption behavior and, ultimately, welfare.
Given the substantial impact of the learning divide, mitigating it and its associated consequences becomes crucial. Our simulation results suggest that providing free training may be an effective intervention. In a single illustrative trajectory shown in Figure Y1, we provide training at the outset by ensuring that the slow learner uses ChatGPT 300 times prior to Day 0. Under this intervention, the user is no longer trapped: the belief is updated more frequently and converge toward the true representative utility of ChatGPT. This example illustrates how early training can help users avoid the belief trap; accordingly, offering training to slow learners may mitigate the learning divide. Our large-scale simulations (Figure Y2) further show that a training program comprising roughly 200–400 uses of ChatGPT allows slow learners to achieve trapping probabilities comparable to those of fast learners, suggesting that this approach can be effective in practice. Finally, Appendix Y further examines (1) the effects of reading news articles and (2) the effects of potential improvements from more advanced large language models. Overall, the amount of news exposure required to alleviate the generative AI divide appears to be very large relative to the average news consumption in our sample. Likewise, the performance improvements needed for generative AI tools to materially reduce the divide are also substantial. These comparisons suggest that training programs are likely to be the more feasible and cost-effective approach to mitigating the generative AI divide.
9 Discussion and Conclusion
9.1 Summary and Discussion
Leveraging a large scale clickstream dataset, we develop and estimate a Bayesian learning model to examine generative AI divide and to identify the underlying mechanisms. Several key findings align with those from prior research on digital divide. Consistent with existing evidence, male users, younger users, and individuals in occupations with greater exposure to LLM learn about generative AI’s utility more rapidly and obtain higher utility per interaction (McElheran et al. 2024, Robinson and others 2020, Scheerder et al. 2017, Dewan and Riggins 2005, Van Dijk and Hacker 2003). Establishing that these patterns persist in the generative AI context is theoretically important, as it shows that familiar forms of inequality remain even when interaction is mediated by advanced AI systems.
More importantly, our results reveal that, conditional on use, users with some college education, users from marginalized racial groups (i.e., non-white users), and users with higher English literacy obtain lower per-use utility. One possible explanation is that generative AI substitutes for users’ baseline cognitive, linguistic, and informational skills rather than complementing them (Noy and Zhang 2023, Haslberger et al. 2023, Autor et al. 1998). Because users with some college education and stronger linguistic or cultural fit possess higher baseline capabilities, they can extract utility in both search engine and generative AI settings, leaving less space for incremental gains from synthesized AI outputs (Agrawal et al. 2024, Humlum and Vestergaard 2024). By contrast, groups that appear disadvantaged along traditional dimensions may realize higher per-use utility because AI-generated synthesized responses substantially reduce search, filtering, and integration costs relative to the search engine environment (Burtch and Chan 2019, Scheerder et al. 2017, Hargittai 2002). This interpretation is consistent with emerging evidence that generative AI can disproportionately benefit lower-skilled or less experienced users by substituting for baseline capabilities, while delivering smaller marginal gains for higher-skilled users who have stronger outside options (Humlum and Vestergaard 2024, Yan et al. 2024, Noy and Zhang 2023). This asymmetric value creation generates a divergence between learning speed and realized utility per-use: some socially advantaged users learn system capabilities more quickly yet obtain smaller benefits per interaction (Vassilakopoulou and Hustad 2023, Lythreatis et al. 2022).
9.2 Contributions and Implications
This study contributes to the literature on the digital divide, AI adoption, and Bayesian learning in several ways. First, we develop a new framework for understanding AI adoption in environments characterized by informational uncertainty and dynamic learning. Whereas prior work on the digital divide often emphasizes static access and initial use, our approach highlights the ongoing learning dynamics that shape how users interact with AI technologies over time. Leveraging real-world behavioral data and a structural Bayesian learning model, we identify self-reinforcing cycles such as the belief trap and disentangle heterogeneity in belief updating from heterogeneity in per-use utility, a distinction that is difficult to recover from survey- or experiment-based designs. Second, we advance the literature on the digital divide by introducing the concepts of the learning divide, utility divide, and belief trap to clarify the mechanisms behind differential AI adoption and outcomes. In doing so, we extend the digital divide frameworks to incorporate dynamic, path-dependent inequalities arising from learning and belief updating under uncertainty. Third, we show that AI technologies, absent appropriate interventions, may amplify existing social inequalities, underscoring the importance of inclusive design and deployment practices.
Beyond these theoretical contributions, our findings also have practical implications for governments, Non-profit organizations (NGOs), AI model providers, and firms. Governments and NGOs can strengthen users’ capabilities to use AI through educational programs that help underprivileged groups overcome information asymmetry. There is an opportunity to go beyond basic access by offering dynamic “on-ramping” supports that reduce signal noise for slow learners. They can consider providing disadvantaged groups with subsidized access to premium, high-fidelity models to reduce the frequency of hallucinations, alongside localized benchmarking tools that help non-native speakers evaluate unstructured outputs more effectively. AI model platforms could also support adoption by facilitating third party evaluations that assess performance across diverse, general-purpose tasks, making system capabilities easier to learn and compare. For organizations providing AI applications to their employees, they have the option to formalize “AI Navigator” peer networks where experienced users share prompt-to-output pairs, transforming private learning into social signals that lower uncertainty for others. Firms that provide AI or embed AI into their products can prioritize inclusive design and offer training programs targeted at disadvantaged users. By addressing the utility divide, learning divide, and belief trap, these stakeholders can promote AI adoption and help ensure that the benefits of AI technologies are broadly shared across society.
9.3 Limitations and Future Work
Our work also highlights several promising directions for future research. First, as in many studies based on clickstream data or online surveys, the social characteristics in our dataset are self-reported and measured with limited granularity (e.g., a binary education indicator). Although the company seeks to balance attributes during panel recruitment, the resulting panel may not perfectly match the U.S. population. In addition, the company cannot guarantee that panelists do not use untracked devices. Since we restrict the sample to individuals with sufficiently high activity levels (at least 60 active days during the observation window), this concern may be partially mitigated, but not eliminated. Despite these data constraints, we hope our study provides a “half answer” to a “big question” and motivates further work on generative AI adoption and digital divide issues (Hosanagar 2017). Future research could apply our framework in other generative AI adoption contexts (e.g., within enterprises) to assess the generalizability of our findings and identify important patterns that vary across scenarios.
Second, generative AI products can undergo updates, and existing Bayesian learning approaches typically do not directly accommodate such non-stationarity (Wang and Yang 2024, Erdem and Keane 1996). In this study, given the currently available toolkit, we conduct robustness checks that focus on the period before the launch of GPT-4 to reduce concerns about version changes (Appendix S). Nonetheless, it would be highly valuable for future methodological work to develop models that explicitly incorporate user learning in the presence of product upgrades, and for future empirical research to leverage such tools to study adoption dynamics more directly.